OpenStudy (anonymous):

What is the equation, in standard form, of a parabola that models the values in the table? x= -2, 0, 4 f(x)= 0, -6, 78

OpenStudy (anonymous):

f(x) is the same meaning as y.

OpenStudy (anonymous):

yes and it is called function of x

OpenStudy (anonymous):

I think you start out f(-2)=4a+(-2b)+3=0 f(0)=ax^2+bx+c=-6 c=-6 f(4)=16a+4b+3=78

OpenStudy (anonymous):

Yes, sortof. You'll need to solve a system of two linear equations.

OpenStudy (anonymous):

You also have x- and y-intercepts, which makes it much easier.

OpenStudy (anonymous):

Start with the standard form template: f(x)=ax^2+bx+c Then put in your known values: f(-2) = a(-2)^2+b(-2)+c = 0 f(0) = a(0)^2+b(0)+c = -6 f(4) = a(4)^2+b(4)+c = 78 From the second equation, you can see that c = -6, so you can substitute that back in the other two.

OpenStudy (anonymous):

You have the right idea, just be careful about how you write the equations.